How the post office boxes were used in post office for finding where the wire breaks? Explain it by drawing the circuit.

Grade:11

2 Answers

Umakant biswal
7 years ago
@ nitesh 
Your query are not clear to me , 
see , post office boxes are besically referred to the wheatstone bridge arrengement . 
in these wheatstone bridge , there are 4 resisters connected in the 4 arms and a galvanometer in the middle arm . 
its used to calculate the unknown resistance used their . 
the formula for calculating the wheatstone bridge unknown resistance .is 
R1/R2 = R3/R4 
u can take in any convenience mode , just the ratio of two resistance will be equal to the ratio of anathor two reisistance . 
and its not used anywhere to check the wire breaks in post officxe . 
KINDLY REPOST THE QUESTION CLEARLY . 
WAITING TO HEAR FROM YOU AGAIN 
jyoti bhatia
7 years ago
How the post office boxes were used in post office for finding where the wire breaks? Explain it by drawing the circuit.
The Post Office Box was a Wheatstone bridge-style testing device with pegs and spring arms to close electrical circuits and measure properties of circuit under test.
The boxes were used in United Kingdom by engineers from the then General Post office, who were responsible for UK telecommunications to trace electrical faults, i.e. to determine where a break occurred in a cable which could be several miles in length. It works on principle of Wheatstone bridge to identify resistance of wire connected and then by using wire resistivity & cross section calculating length of wire connected & thus determining where cable had broken.
 A typical Post Office Box is in a wooden box with a hinged lid and a metal or bakelite panel showing circuit connections. Coils of wire are wound non-inductively, mounted in body of the box and have a negligible temperature coefficient.
Pairs of ratio arms are each 10, 100, 1000 ohms. Resistance arms contain a number of coils from 1 to 5000 ohms with a plug for infinite resistance.
 
Wheatstone bridge circuit diagram: Unknown resistance Rx is to be measured; resistances R1, R2 and R3 are known and R2 is adjustable. If measured voltage Vg is zero, then R2/R1 = Rx/R3.
A Wheatstone bridge is an electrical circuit used to measure an unknown electrical resistance by balancing two legs of a bridge circuit, one leg of which includes unknown component. Primary benefit of a Wheatstone bridge is its ability to provide extremely measurements. Its operation is similar to the original potentiometer.